OpenStudy (anonymous):

In which quadrant does the terminal side of a 117° angle lie?

OpenStudy (anonymous):

Think that \[ 90 < 117 < 180 \] and the first quadrant is \[0 \le \theta \le 90 \] What is your answer now?

OpenStudy (anonymous):

2nd quadrant